# A randomized lattice rule without component-by-component construction

### v1

@@ 4.1. Distribution of the worst-case error
First, we empirically investigate the distribution of the worst-case error in the same weighted Korobov space using two different randomized lattice rules: our new algorithm (Algorithm 3.1) and the randomized CBC construction introduced by [5]. Our new algorithm selects a generating vector by randomly drawing generating vectors multiple times from all possible candidates and choosing the one that minimizes the worst-case error among them. This ensures that the worst-case error achieves nearly the optimal order with a very high probability. On the other hand, any generating vector obtained from the randomized CBC construction guarantees that the worst-case error of the corresponding lattice rule achieves nearly the optimal order with probability 1. Here, we aim to compare the worst-case error distributions of rank-1 lattice rules obtained from these two algorithms.
@@ 4.2. Performance for test functions
with $\beta=2,3,4$. It can be verified that $f_{1}\in H_{2,\boldsymbol{\gamma}}$ but $f_{1}\not\in H_{3,\boldsymbol{\gamma}}$. Similarly, $f_{\beta}\in H_{\beta,\boldsymbol{\gamma}}$ but $f_{\beta}\not\in H_{\beta+1,\boldsymbol{\gamma}}$ for any $\beta=2,3,4$. Note that the true integral for all these test functions is equal to $1$.
We present the results for all the test functions in the lower-dimensional case of $d=2$ in Figure 2. The horizontal axis represents the maximum number of points $M$, while the vertical axis displays the sample variance (both on a base-$10$ logarithmic scale). Additionally, we include the results for the standard Monte Carlo method with $M$ random points as a reference. As anticipated, the variance of the standard Monte Carlo method decreases at a rate of $O(M^{-1})$. In contrast, the two randomized lattice rules demonstrate much faster convergence behaviors, achieving a variance rate of $O(M^{-5})$ for $f_{1}$ and even better rates for other functions. Due to machine precision limitations, the variance no longer decreases beyond a certain point for $f_{3}$ and $f_{4}$. This observation agrees well with our theoretical findings. In this setting, our new randomized rule exhibits slightly better performance than the randomized CBC approach overall.
As illustrated in Figure 3, high-order convergence behaviors of both the randomized lattice rules are maintained even in the high-dimensional setting with $d=20$. Although the randomized CBC approach shows slightly superior performance for $f_{2}$ compared to our new algorithm, the performance of both approaches is quite comparable for other functions. Hence, in the randomized setting, lattice rules based on the CBC construction are not the only viable choice, suggesting potential for the development of alternative implementable approaches with both theoretical and empirical support.

### v3

@@ 4.1. Distribution of the worst-case error
First, we empirically investigate the distribution of the worst-case error in the same weighted Korobov space using two different randomized lattice rules: our new algorithm (Algorithm 3.1) and the randomized CBC construction introduced by [5]. Our new algorithm selects a generating vector by randomly drawing generating vectors multiple times from all possible candidates and choosing the one that minimizes the worst-case error among them. This ensures that the worst-case error achieves nearly the optimal order with a very high probability. In fact, as shown in Lemma 3.2, the probability of the event $\boldsymbol{z}^{*}\in\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}$ is at least $1-(1-\eta)^{r}$. On the other hand, any generating vector obtained from the randomized CBC construction guarantees that the worst-case error of the corresponding lattice rule achieves nearly the optimal order with probability 1. Here, we aim to compare the worst-case error distributions of rank-1 lattice rules obtained from these two algorithms.
@@ 4.2. Performance for test functions
with $\beta=2,3,4$. It can be verified that $f_{1}\in H_{5/2-\varepsilon,\boldsymbol{\gamma}}$ and $f_{\beta}\in H_{(2\beta+1)/2-\varepsilon,\boldsymbol{\gamma}}$ for any $\beta=2,3,4$ with arbitrarily small $\varepsilon>0$. This implies that the best possible convergence rates of the variance are $M^{-6+\varepsilon},M^{-6+\varepsilon},M^{-8+\varepsilon},M^{-10+\varepsilon}$, respectively. Note that the true integral for all these test functions is equal to $1$.
We present the results for all the test functions in the lower-dimensional case of $d=2$ in Figure 2. The horizontal axis represents the maximum number of points $M$, while the vertical axis displays the sample variance (both on a base-$10$ logarithmic scale). Additionally, we include the results for the standard Monte Carlo method with $M$ random points as a reference. As anticipated, the variance of the standard Monte Carlo method decreases at a rate of $O(M^{-1})$. In contrast, the two randomized lattice rules demonstrate much faster convergence behaviors. Due to machine precision limitations, the variance no longer decreases beyond a certain point for $f_{3}$ and $f_{4}$. Also, we see that our new randomized rule exhibits slightly better performance than the randomized CBC approach overall. The linear regression for the results in the suitable ranges of $M$ indicates that our proposed randomized lattice rule achieves variance rates of $M^{-5.67},M^{-7.30},M^{-7.04},M^{-9.16}$, respectively. This observation almost agrees with our theoretical findings. It is noteworthy that while we theoretically expect better convergence for $f_{3}$ compared to $f_{2}$, the experiment shows the opposite result. Investigating the reason for this discrepancy is beyond the scope of this work.
As illustrated in Figure 3, high-order convergence behaviors of both the randomized lattice rules are maintained even in the high-dimensional setting with $d=20$. However, the rate of convergence deteriorates considerably; the linear regression suggests that our proposed randomized lattice rule achieves variance rates of $M^{-4.21},M^{-4.09},M^{-5.46},M^{-6.77}$, respectively. Although we omit the details here, the rate of convergence can be significantly improved, for example, by replacing the weights $1/j^{2\beta}$ in $f_{\beta}$ for $\beta=2,3,4$ with faster-decaying weights. This confirms the importance of weight parameters in high-dimensional settings. In our current experiments, although the randomized CBC approach shows slightly superior performance for $f_{2}$ compared to our new algorithm, the performance of both approaches is quite comparable for other functions. Hence, in the randomized setting, lattice rules based on the CBC construction are not the only viable choice, suggesting potential for the development of alternative implementable approaches with both theoretical and empirical support.
Finally, we briefly compare our results with those of a deterministic CBC-constructed rank-1 lattice rule. Numerical experiments for the same test functions showed that, in most cases, the (squared) absolute error decay rate of the deterministic rule was comparable to, or even better than, the variance decay rate achieved by the randomized rank-1 lattice rules. This can be attributed to the following: considering a Banach space with the norm
$\|f\|_{\alpha,\boldsymbol{\gamma}}:=\sup_{\boldsymbol{k}\in\mathbb{Z}^{d}}|\hat{f}(\boldsymbol{k})|\,r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}),$
our test functions $f_{1},\ldots,f_{4}$ belong to this space with smoothness parameters $\alpha=3,3,4,5$, respectively. In such spaces, a deterministic rank-1 lattice rule can achieve a worst-case error of order $N^{-\alpha+\varepsilon}$ for arbitrarily small $\varepsilon>0$. Moreover, improving this rate by considering randomized error criteria and randomized rank-1 lattice rules seems quite difficult. Therefore, determining the superiority of deterministic versus randomized algorithms for specific functions, in terms of their convergence rates, is not straightforward and requires further investigation.